The collision impact of an automobile varies jointly as its mass and the square of it's speed. Suppose a 2000-lb car traveling at 55mph has a collision impact of 12.3. What is the collision impact of the same car at 65mph?
Answer by stanbon(75887)

The collision impact of an automobile varies jointly as its mass and the square of it's speed.
i = k*m*s^2
---
Solve for "k" using "a 2000-lb car traveling at 55mph has a collision impact of 12.3".
12.3 = k*2000*55^2
k = 12.3/(2000*55^2)
k = 0.000002033
------------------------
Equation:
i = 0.000002033*m*s^2
==============================
What is the collision impact of the same car at 65mph?
i = 0.000002033*2000*65^2 = 17.18
==============
Cheers,
Stan H.